# Task-Agnostic Machine Learning-Assisted Inference

### v1

@@ 2.3 Related work
Our work is closely related to recent methods developed in the ML-assisted inference literature (Angelopoulos et al., 2023a, ; Miao et al., 2023b, ; Wang et al.,, 2020; Angelopoulos et al., 2023b, ; Gan and Liang,, 2023; Motwani and Witten,, 2023; Zrnic and Candès,, 2024), and is also related to methods for handling missing data (Rubin,, 1996; Tsiatis,, 2006) and semi-supervised inference (Azriel et al.,, 2022; Zhang and Bradic,, 2022; Deng et al.,, 2023; Zhang et al.,, 2019). While current ML-assisted inference methods modify the loss function or the estimating equation, our protocol works directly on the summary statistics. For simple problems such as mean estimation, current methods yield a closed-form solution to the optimization problem. However, for more general statistical tasks, there is no such closed-form solution. Current methods typically require the algebraic form of the loss function, its first and second order derivatives, and the standard error for the estimator, as well as a newly implemented optimization algorithm to obtain the estimator. We use the logistic regression problem as an example. Here, ${\bm{\theta}}^{*}=\mbox{arg min}_{{\bm{\theta}}}\mathbb{E}[-Y({\bm{\theta}}{\bf X})^{\rm T}-\psi({\bf X}{\bm{\theta}})]$ and $\psi(t)=1/(1+\exp(-t))$.
[…]
with asymptotic variance ${\bf A}^{-1}{\bf V}(\omega){\bf A}^{-1}$, where ${\bf A}=\frac{1}{N+n}(\sum_{i=1}^{n}\psi^{\prime\prime}\left({\bf X}_{i}{\bm{\theta}}\right){\bf X}_{i}^{\rm T}{\bf X}_{i}+\sum_{j=1}^{N}\psi^{\prime\prime}\left({\bf X}_{i}{\bm{\theta}}\right){\bf X}_{i}^{\rm T}{\bf X}_{i})$, ${\bf V}(\omega)=\frac{n}{N}[\omega^{2}\operatorname{Var}_{n}\left((\psi^{\prime}({\bf X}_{i}{\bm{\theta}})-\widehat{f}_{i}){\bf X}_{i}^{\rm T}\right)+\operatorname{Cov}_{N+n}\left((1-\omega)\psi^{\prime}({\bf X}_{i}{\bm{\theta}}){\bf X}_{i}^{\rm T}+(\omega\widehat{f}_{i}-Y_{i}){\bf X}_{i}^{\rm T}\right)$, and $\omega$ needs to be obtained by optimization to minimize the asymptotic variance. In contrast, our protocol simply applies logistic regression $\mathcal{A}$ to
@@
  • labeled data with observed outcomes $({\bf X}_{\cal L},Y_{\cal L})$ to obtain $[\widehat{\bm{\theta}}_{{\cal L}},\operatorname{Var}(\widehat{\bm{\theta}}_{{\cal L}})]$
@@
  • labeled data with predicted outcomes $({\bf X}_{\cal L},\widehat{f}_{\cal L})$ to obtain [$\widehat{\bm{\eta}}_{{\cal L}},\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal L}})]$
@@
  • unlabeled data with predicted outcomes $({\bf X}_{\cal U},\widehat{f}_{\cal U})$ to obtain $[\widehat{\bm{\eta}}_{{\cal U}},\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal U}})]$
  • bootstrap of labeled data $({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q$ for $\operatorname{Cov}(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}})$,
@@
and returns $\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{\bm{\theta}}_{{\cal L}})$, where $\bm{\omega}_{0}=(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$. For each new statistical task, as long as an existing analysis routine can produce an estimator that is asymptotically normally distributed, our protocol can be applied in a similar manner. Additionally, the current mathematical principles guiding ML-assisted inference are applicable solely to M-estimation (Angelopoulos et al., 2023a, ; Miao et al., 2023b, ; Angelopoulos et al., 2023b, ; Gan and Liang,, 2023; Zrnic and Candès,, 2024). However, our protocol extends beyond this limitation, addressing all estimation problems with an asymptotically normally distributed estimator.
@@
Inference relying solely on summary statistics is widely used in the statistical genetics literature for practical reasons. Summary statistics-based methods have been developed for tasks such as variance component inference and polygenic risk prediction (Bulik-Sullivan et al., 2015b, ; Bulik-Sullivan et al., 2015a, ; Ruan et al.,, 2022; Miao et al., 2023a, ; Zhao et al.,, 2022). In contrast to our work, these applications do not leverage ML predictions, but instead focus on inference using summary statistics obtained from observed outcomes. An exception is our previous study for valid genome-wide association studies on ML-predicted outcome (Miao et al.,, 2024). But our previous work focused only on linear regression modeling with application to genome-wide association studies. The PSPS introduced in this paper aims to extend ML-assisted inference to general statistical tasks.
@@
Our work is also related to resampling-based inference, zero augmentation, and false discovery rate (FDR) control methods. Our protocol is inspired by the core principle of resampling-based inference, which replaces algebraic derivation with computation (Efron and Tibshirani,, 1994). The main difference is that we focus on how to use ML to support inference, whereas resampling-based inference focuses on bias and variance estimation. The idea of zero augmentation has been used in augmented inverse propensity weighting estimators (Robins et al.,, 1994) and in handling unmeasured confounders (Yang and Ding,, 2019) and missing data for U-statistics (Cannings and Fan,, 2022). These estimators do not incorporate ML, which is fundamental to our work. We also adapt techniques from the FDR literature (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001; Barber and Candès,, 2015, 2019). Our unique contribution is to use ML to support FDR control, thereby increasing its statistical power, in contrast to classical methods that rely solely on labeled data.
@@ 3.1 General protocol for PSPS
Building on our two examples in section 2, we formalize our protocol in fig. 1 for ML-assisted inference:
Algorithm 1 PSPS for ML-assisted inference
@@
  • labeled data $({\bf X}_{\cal L},Y_{\cal L})$ and obtain $[\widehat{\bm{\theta}}_{\cal L},\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})]$
  • labeled data $({\bf X}_{\cal L},\widehat{f}_{\cal L})$ and obtain $[\widehat{\bm{\eta}}_{\cal L},\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})]$
  • unlabeled data with $({\bf X}_{\cal U},\widehat{f}_{\cal U})$ and obtain $[\widehat{\bm{\eta}}_{\cal U},\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U})]$
  • $Q$ bootstrap of labeled data $({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q$ and obtain $\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$.
4:Employ one-step debiasing to the the summary statistics in step2: $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{{\bm{\theta}}}_{{\cal L}}),$ where $\bm{\omega}_{0}=(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ and $\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})=\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})-\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}(\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})+\frac{n}{N}\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$
5:ML-assisted point estimator, standard error, confidence interval, and p-value
The only requirement for our protocol are: i) algorithm $\mathcal{A}$, when applied to labeled data $(\bm{X}_{\cal L},Y_{\cal L})$, returns a consistent and asymptotically normally distributed estimator of ${\bm{\theta}}^{*}$; ii) labeled and unlabeled data are i.i.d.. Under these assumptions, the summary statistics asymptotically follows:
@@
where $\bm{\eta}\equiv\bm{\eta}(\mathbb{P}_{\widehat{f}})\in\mathbb{R}^{K}$ is defined on $({\bf X},\widehat{f})\sim\mathbb{P}_{\widehat{f}}$, ${\bf V}(\cdot)$ denotes the asymptotic variance and covariance of a estimator, and $\rho=\frac{n}{N}$. The asymptotic approximation gives ${\bf V}(\widehat{\bm{\theta}}_{\cal L})=n\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L}),{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})=n\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L}),{\bf V}(\widehat{\bm{\eta}}_{\cal L})=n\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})$ and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})=N\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U})$. Here, we do not require $\widehat{\bm{\eta}}_{\cal L}$ and $\widehat{\bm{\eta}}_{\cal U}$ to be consistent for $\bm{\theta}^{*}$, thus allows arbitrary ML model.
@@
With the asymptotically multivariate normal distribution for the summary statistics, the one-step debiased estimator $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{{\bm{\theta}}}_{{\cal L}})$ is consistent for $\theta^{*}$ and asymptotically normally distributed. The $\alpha$-level confidence interval for coordinate $k$ can be constructed by $\mathcal{C}_{\alpha}^{\textnormal{{PSPS}}}=(\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}\pm z_{1-\alpha/2}\sqrt{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})$. Hypothesis testing for $H_{0}:\theta^{*}_{k}=0$ can be performed using the test statistics $\frac{\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}}{\sqrt{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})}\sim\mathcal{N}(0,1)$. It gives (two-sided) p-value $=2(1-\Phi(|\frac{\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}}{\sqrt{\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})}|))$, where $\Phi$ is the CDF of the standard normal distribution.

### v3

@@ 2.3 Related work
Our work is closely related to recent methods developed in the literature of ML-assisted inference [3, 35, 46, 4, 20, 37, 56], and is also related to methods for handling missing data [40, 42] and semi-supervised inference [6, 52, 16, 50]. While current ML-assisted inference methods modify the loss function or the estimating equation, our protocol works directly on the summary statistics. For simple problems such as mean estimation, current methods yield a closed-form solution to the optimization problem. However, for more general statistical tasks, there is no such closed-form solution. Current methods typically require the algebraic form of the loss function, its first- and second-order derivatives, and the variance for the estimator, as well as a newly implemented optimization algorithm to obtain the estimator. We use the logistic regression problem as an example. Here, ${\bm{\theta}}^{*}=\mbox{arg min}_{{\bm{\theta}}}\mathbb{E}[-Y({\bm{\theta}}{\bf X})^{\rm T}-\psi({\bf X}{\bm{\theta}})]$ and $\psi(t)=1/(1+\exp(-t))$. The ML-assisted estimator is $\widehat{\theta}_{\texttt{MLA}}=\mbox{arg min}_{{\bm{\theta}}}\frac{1}{N}\sum\limits_{i=n+1}^{n+N}\widehat{\omega}[-\widehat{f}_{i}{\bm{\theta}}^{\rm T}{\bf X}_{i}^{\rm T}-\psi({\bf X}_{i}{\bm{\theta}})]-\{\frac{1}{n}\sum\limits_{i=1}^{n}\widehat{\omega}[-\widehat{f}_{i}{\bm{\theta}}^{\rm T}{\bf X}_{i}^{\rm T}-\psi({\bf X}_{i}{\bm{\theta}})]-\frac{1}{n}\sum\limits_{i=1}^{n}[-\widehat{Y}_{i}{\bm{\theta}}^{\rm T}{\bf X}_{i}^{\rm T}-\psi({\bf X}_{i}{\bm{\theta}})]\}$ with estimated asymptotic variance $\widehat{\bf A}^{-1}\widehat{\bf V}(\omega)\widehat{\bf A}^{-1}$, where $\widehat{\bf A}=\frac{1}{N+n}(\sum_{i=1}^{n}\psi^{\prime\prime}\left({\bf X}_{i}{\bm{\theta}}\right){\bf X}_{i}^{\rm T}{\bf X}_{i}+\sum_{i=n+1}^{n+N}\psi^{\prime\prime}\left({\bf X}_{i}{\bm{\theta}}\right){\bf X}_{i}^{\rm T}{\bf X}_{i})$, $\widehat{\bf V}(\omega)=\frac{n}{N}[\widehat{\omega}^{2}\operatorname{Var}_{n}\left((\psi^{\prime}({\bf X}_{i}{\bm{\theta}})-\widehat{f}_{i}){\bf X}_{i}^{\rm T}\right)+\widehat{\operatorname{Cov}}_{N+n}\left((1-\widehat{\omega})\psi^{\prime}({\bf X}_{i}{\bm{\theta}}){\bf X}_{i}^{\rm T}+(\widehat{\omega}\widehat{f}_{i}-Y_{i}){\bf X}_{i}^{\rm T}\right)$, and $\widehat{\omega}$ needs to be obtained by optimization to minimize the asymptotic variance. In contrast, our protocol simply applies logistic regression $\mathcal{A}$ to
@@
  • labeled data with observed outcomes $({\bf X}_{\cal L},Y_{\cal L})$ to obtain $[\widehat{\bm{\theta}}_{{\cal L}},\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{{\cal L}})]$
@@
  • labeled data with predicted outcomes $({\bf X}_{\cal L},\widehat{f}_{\cal L})$ to obtain [$\widehat{\bm{\eta}}_{{\cal L}},\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{{\cal L}})]$
@@
  • unlabeled data with predicted outcomes $({\bf X}_{\cal U},\widehat{f}_{\cal U})$ to obtain $[\widehat{\bm{\eta}}_{{\cal U}},\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{{\cal U}})]$
  • bootstrap of labeled data $({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q$ for $\widehat{\operatorname{Cov}}(\widehat{\bm{\theta}}_{{\cal L}},\widehat{\bm{\eta}}_{{\cal L}})$,
@@
and returns $\widehat{\bm{\omega}}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{\bm{\theta}}_{{\cal L}})$, where $\widehat{\bm{\omega}}_{0}=(\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal L})+\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal U}))^{-1}\widehat{\operatorname{Cov}}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$. For each new statistical task, as long as an existing analysis routine can produce an estimator that is asymptotically normally distributed, our protocol can be similarly applied. Additionally, the current mathematical principles guiding ML-assisted inference apply solely to M-estimation [3, 35, 4, 20, 56]. Our protocol extends beyond this limitation, addressing all estimation problems with an asymptotically normally distributed estimator.
@@
Inference relying solely on summary statistics is widely used in the statistical genetics literature for practical reasons. Summary statistics-based methods have been developed for tasks such as variance component inference and genetic risk prediction [12, 11, 39, 34, 53]. In contrast to our work, these applications do not leverage ML predictions, but instead focus on inference using summary statistics obtained from observed outcomes. An exception is a previous study for valid genome-wide association studies (GWAS) on ML-predicted outcome [36]. However, it focused only on linear regression modeling with application to GWAS. The PSPS  framework introduced in this paper aims to extend ML-assisted inference to general statistical tasks.
@@
Our work is also related to semi-supervised learning, resampling-based inference, zero augmentation, and false discovery rate (FDR) control methods. Our protocol is designed for estimation and statistical inference using both labeled and unlabeled data, addressing a different problem from semi-supervised learning [55], which primarily focuses on prediction. Our protocol is inspired by the core principle of resampling-based inference, which replaces algebraic derivation with computation [19]. The main difference is that we focus on how to use ML to support inference, whereas resampling-based inference focuses on bias and variance estimation, and type-I error control. The idea of zero augmentation has been used in augmented inverse propensity weighting estimators [38] and in handling unmeasured confounders [48] and missing data for U-statistics [14]. These estimators do not incorporate ML, which is fundamental to our work. We also adapt techniques from the FDR literature [9, 10, 7, 8]. Our unique contribution is to use ML to support FDR control, thereby increasing its statistical power, in contrast to classical methods that rely solely on labeled data.
@@ 3.1 General protocol for PSPS
Building on section 2, we formalized our protocol in fig. 1 for ML-assisted inference:
@@
  • labeled data $({\bf X}_{\cal L},Y_{\cal L})$ and obtain $[\widehat{\bm{\theta}}_{\cal L},\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\cal L})]$
  • labeled data $({\bf X}_{\cal L},\widehat{f}_{\cal L})$ and obtain $[\widehat{\bm{\eta}}_{\cal L},\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal L})]$
  • unlabeled data with $({\bf X}_{\cal U},\widehat{f}_{\cal U})$ and obtain $[\widehat{\bm{\eta}}_{\cal U},\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal U})]$
  • $Q$ bootstrap of labeled data $({\bf X}_{\cal L},Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q$ and obtain $\widehat{\operatorname{Cov}}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$.
4:Employ one-step debiasing to the summary statistics in step2: $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\widehat{\bm{\omega}}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\widehat{\bm{\omega}}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{{\bm{\theta}}}_{{\cal L}}),$ where $\widehat{\bm{\omega}}_{0}=[\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal L})+\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal U})]^{-1}\widehat{\operatorname{Cov}}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ and $\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})=\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\cal L})-\widehat{\operatorname{Cov}}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}[\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal L})+\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal U})]^{-1}\widehat{\operatorname{Cov}}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$
5:ML-assisted point estimator $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}$, standard error $\sqrt{\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})}$, $\alpha$-level confidence interval for the $k$-th coordinate $\mathcal{C}_{\alpha,k}^{\textnormal{{PSPS}}}=(\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}\pm z_{1-\alpha/2}\sqrt{\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})$, and (two-sided) p-value $2(1-\Phi(|\frac{\widehat{{\bm{\theta}}}_{\textnormal{{PSPS}}_{k}}}{\sqrt{\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}})}|))$, where $\Phi$ is the CDF of the standard normal distribution.
Algorithm 1 PSPS for ML-assisted inference
The only requirements for our protocol are: i) algorithm $\mathcal{A}$, when applied to labeled data $(\bm{X}_{\cal L},Y_{\cal L})$, returns a consistent and asymptotically normally distributed estimator of ${\bm{\theta}}^{*}$; ii) labeled and unlabeled data are independent and identically distributed. Under these assumptions, the summary statistics have the following asymptotic properties:
@@
where $\bm{\eta}\equiv\bm{\eta}(\mathbb{P}_{\widehat{f}})\in\mathbb{R}^{K}$ is defined on $({\bf X},\widehat{f})\sim\mathbb{P}_{\widehat{f}}$, ${\bf V}(\cdot)$ denotes the asymptotic variance and covariance of a estimator, and $\rho=\frac{n}{N}$. The asymptotic approximation gives ${\bf V}(\widehat{\bm{\theta}}_{\cal L})\approx n\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L}),{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})\approx n\operatorname{Cov}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L}),{\bf V}(\widehat{\bm{\eta}}_{\cal L})\approx n\operatorname{Var}(\widehat{\bm{\eta}}_{\cal L})$ and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})\approx N\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U})$. Here, we do not require $\widehat{\bm{\eta}}_{\cal L}$ and $\widehat{\bm{\eta}}_{\cal U}$ to be consistent for ${\bm{\theta}}^{*}$, thus allows arbitrary ML model.
@@
With the summary statistics following a multivariate normal distribution asymptotically, the debiased estimator $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}=\widehat{\bm{\omega}}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\widehat{\bm{\omega}}_{0}^{\rm T}\widehat{\bm{\eta}}_{{\cal L}}-\widehat{{\bm{\theta}}}_{{\cal L}})$ is consistent for $\theta^{*}$ and asymptotically normally distributed (theorem 1). Therefore, by plugging in a consistent estimator for its asymptotic variance ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})\approx n\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})$, valid confidence interval and hypothesis testing can be achieved.
Remark 1.
PSPS is more "task-agnostic" than existing methods in three aspects:
  1. 1.
    For M-estimation tasks, currently, only mean and quantile estimation, as well as linear, logistic, and Poisson regression, have been implemented in software tools and are ready for immediate application. For other M-estimation tasks, task-specific derivation of the ML-assisted loss functions and asymptotic variance via the central limit theorem are necessary. After that, researchers still need to develop software packages and optimization algorithms to carry out real applications. In contrast, PSPS only requires already implemented algorithms and software designed for classical inference based on labeled data.
  2. 2.
    For problems that do not fall under M-estimation but have asymptotically normally distributed estimators, only PSPS can be applied, and all current methods would fail. The principles behind ML-assisted M-estimation do not extend to these tasks.
  3. 3.
    Even for M-estimation tasks that have already been implemented, PSPS offers the additional advantage of relying solely on summary statistics. The “task-specific derivations” refer not only to statistical tasks but also to scientific tasks. Real-world data analysis in any scientific discipline often involves conventions and nuisances that require careful consideration. For example, our work is partly motivated by GWAS [43]. Statistically, GWAS is a linear regression that regresses an outcome on many genetic variants. While the regression-based statistical foundation is simple, conducting a valid GWAS requires accounting for numerous technical issues, such as sample relatedness (i.e., study participants may be genetically related) and population structure (i.e., unrelated individuals of the same ancestry are both genetically and phenotypically similar, creating confounded associations in GWAS). Sophisticated algorithms and software have been developed to address these complex issues [31]. It will be very challenging if all these important features need to be reimplemented in an ML-assisted GWAS framework. With our PSPS protocol, researchers can utilize existing tools that are highly optimized for genetic applications to perform ML-assisted GWAS. This adaptability is not just limited to GWAS, but is a major feature of our approach across scientific domains. PSPS enables researchers to conduct ML-assisted inference using well-established data analysis routines.
Remark 2.
The "federated data analysis" feature of PSPS refers to the fact that we only require summary statistics as input for inference, rather than individual-level raw data (features ${\bf X}$ and label $Y$). For example, consider a scenario where labeled data is in one center and unlabeled data is in another, yet researchers cannot access individual-level data from both centers simultaneously. Under such conditions, current ML-assisted inference, which relies on accessing both labeled and unlabeled data to minimize a joint loss function, is not feasible. However, PSPS circumvents this issue by aggregating summary statistics from multiple centers, thereby performing statistical inference while upholding the privacy of individual-level data.